OpenStudy (anonymous):

If someone could walk me through the steps on how to solve this I'll fan and medal. Solve the following system of equations by using the elimination method. x – y = 11 2x + y = 19 A. (1,10) B. (-1,-1) C. (12,2) D. (10,-1)

OpenStudy (anonymous):

I will try to solve your system of equations. x−y=11;2x+y=19 Step: Solve x−y=11 for x: x−y+y=11+y(Add y to both sides) x=y+11 Step: Substitute (y+11) for x in 2x+y=19: 2x+y=19 2(y+11)+y=19 3y+22=19(Simplify both sides of the equation) 3y+22+−22=19+−22(Add -22 to both sides) 3y=−3 3y/3 = −3/3 (Divide both sides by 3) y=−1 Step: Substitute (−1) for y in x=y+11: x=y+11 x=−1+11 x=10(Simplify both sides of the equation) Answer: x=10 and y=−1

OpenStudy (anonymous):

Thank you so much for taking the time to help me!

OpenStudy (anonymous):

Your Welcome! :) I hope you understand, if you don't I can always try to help you walk through it by saying the step by step. If you would like.

OpenStudy (anonymous):

I think I understand. There were some things that I didn't understand and I learn by watching how the work is done. Thank you

OpenStudy (anonymous):

Okay. You Welcome!